Calculate the volume of .5 mole of gas at 140K and a pressure of 1.2atm
1
Expert's answer
2020-02-20T07:45:15-0500

It is given that number of moles of gas =0.5 mole

Temperature of gas =140 K

Pressure = 1.2 atm

we know that , using ideal gas equation


pV= nRT

here, R= gas constant whose value is .082 L atm mol -1K -1

"1.2 \times V = 0.5 \times 0.082\times 140"

"V= \frac{0.5\times0.082\times 140}{1.2}"

V=4.783 L

So, volume of gas will be V= 4.783 L